OpenStudy (anonymous):

write the polynomial in factored form x^3-3x^2-10x

OpenStudy (anonymous):

take x as a common factor then we have \[x(x^2 -3x -10)\] now factor the polynomial \[x(x^2 -3x -10) = x(x^2 -5x +2x -10)\] \[x(x^2 -5x +2x -10) = x[x(x-5)+2(x-5)]\] \[x[(x-5)(x+2)]\]

OpenStudy (anonymous):

is that the final answer

OpenStudy (anonymous):

nopes you have to multiply the x to the polynomial in the bracket

OpenStudy (anonymous):

is the final answer 11

OpenStudy (anonymous):

Actually I thought wrong. x[(x-5)(x+2)] is the final ans